the sum of the ages of father and son is 55 years.after 16 years the father will be twice as old as his son, what are their present ages?
Answer by mananth(16946)

let the age of son be x
father's age = 55-x
..
after 16 years their ages will be:
son x+16
father 55-x+16= 71-x
..
Father's age = twice son's age
71-x= 2(x+16)
71-x = 2x +32
3x= 71-32
3x=39
x=13 years , son's age
Father's age = 55-x
= 55-13
=42 years
Your feedback is very important for improving the solution